# Star Attention: Efficient LLM Inference over Long Sequences

### v2

@@ 1 Introduction
The importance of long-context capabilities has driven substantial research into addressing the computational challenges of self-attention. Some approaches focus on reducing the need to fully materialize the attention matrix (Milakov & Gimelshein, 2018), leading to blockwise processing techniques (Dao et al., 2022; Dao, 2024; Liu & Abbeel, 2024) and further optimization through distributed computation across multiple compute units (Liu et al., 2024a). While these methods improve training efficiency, autoregressive decoding during inference still requires the model to attend to every previous token, resulting in high computational costs for long-context sequences. Other approaches attempt to optimize inference by segmenting long inputs into chunks, encoding them separately, and attending to these encoded chunks using the user query (Beltagy et al., 2020; Russak et al., 2024; Liao et al., 2024). However, these methods often require fine-tuning the model or introducing additional components that necessitate further training, limiting their out-of-the-box applicability.
We introduce Star Attention111Code is available at: https://github.com/NVIDIA/Star-Attention, a novel algorithm for efficient LLM long-context inference. This method is based on the observation that LLM inference usually has two stages: (1) prompt encoding, where the model processes input and stores KV vectors in the cache and (2) token generation, where model attends to the KV cache and autoregressively generates new tokens while updating the cache with the new KV vectors. In many long-context tasks, the input consists of a long context followed by a short query and a short answer. The information needed for answering the query is often localized within small parts of the context, meaning context tokens need only attend to nearby tokens, while query tokens need to attend to all prior tokens. Based on this observation, Star Attention utilizes a two-phase approach shown in Figure 1:
@@ 2.1 Phase 1: Context Encoding
where each augmented block contains tokens: tokens from the anchor block followed by tokens from the current block (Figure 2.1). The positional indices of are preserved, ensuring that its tokens retain their original position indices . The augmented blocks are distributed across compute hosts, where each host computes attention over the tokens from its assigned block and generates the corresponding key-value (KV) vectors. While KVs for the anchor block are discarded, the KVs for the current block are retained in the cache.
—c—c—c—c—c—c—c—
[…]
@@ 3 Experiments
Model | Seq. Len. | Block Size | Ring-Attn | Star-Attn | |
(K) | (K) | Acc.(%) | Acc. | Speedup | |
16 | 4 | 92.22 | -0.94% | 1.1x | |
32 | 8 | 87.53 | +1.17% | 1.2x | |
64 | 16 | 84.79 | -1.42% | 1.8x | |
Llama-3.1-8B-Instruct Meta-AI (2024) | 128 | 32 | 76.31 | -1.90% | 2.7x |
16 | 4 | 95.09 | -2.71% | 1.7x | |
32 | 8 | 94.61 | -2.55% | 2.0x | |
Llama-3.1-70B-Instruct Meta-AI (2024) | 64 | 16 | 88.54 | -1.44% | 4.7x |
@@ 3.2 Results
3.3 Comparison with Other Sparse Attention Methods
@@ 3.4 Trade-off between accuracy and speed
Empirically, setting the block size to approximately one-quarter of the total sequence length strikes an effective trade-off between accuracy and speed. For sequence lengths exceeding 128K, we fix the block size at 32K tokens to prioritize inference speed. As shown in Figure 6(a), this configuration allows Star Attention to achieve substantial speedups over Ring Attention while incurring only modest accuracy degradation. For instance, on the RULER benchmark with Llama-3-8B-Instruct-1048K, Star Attention achieves up to 11 speedup while retaining accuracy comparable to Ring Attention. At 1M tokens, the speedup increases to 16.9 with an accuracy drop of just 5.32%.
@@ 3.5 In-Depth Analysis on RULER Task Categories
To better understand the strengths and limitations of Star Attention, we analyze its performance across different task categories within the RULER benchmark. RULER comprises five categories: Single-NIAH, Multi-NIAH, Multi-Hop Tracing, Aggregation, and Question Answering (QA). Figure 6(b) reports category-wise accuracy using the Llama-3.1-8B-Instruct model at a sequence length of 32K and a block size of 8K. We observe consistent trends across all sequence lengths, as detailed in Appendix D.
@@ 4 Ablation Study
randomly sampled from [0, current_block) | 96.79 | -2.71% | 97.16 | -1.33% |
same as previous block | 97.35 | -2.15% | 96.80 | -1.69% |
same as first block | 97.61 | -1.89% | 97.54 | -0.95% |
@@
constant token (ex: ‘ ’ or ‘ the’ or ‘.’ ) | 0.00 | -99.5% | 0 | -98.49% |
random tokens | 90.55 | -8.95% | 82.63 | -15.86% |
shuffled first block tokens | 92.96 | -6.54% | 90.76 | -7.73% |
first block tokens | 97.61 | -1.89% | 94.94 | -3.55% |
Previous-block used as anchor | 94.20 | -5.3% | 96.13 | -2.66% |
@@ 5 Related Work
Memory Optimization: Maintaining the KV cache during autoregressive decoding is a major memory bottleneck. KV cache compression (Ge et al., 2024; Munkhdalai et al., 2024; Sun et al., 2024; Liu et al., 2024b; Wu et al., 2024) and low-rank approximation methods (Srebro & Jaakkola, 2003) have been proposed to trade precision for reduced memory. Recent systems explore eviction-based memory management strategies that allow LLMs to operate over virtually infinite contexts (Zhao et al., 2024; Han et al., 2024; Xiao et al., 2024a), often requiring architecture changes or specialized runtime support. Star Attention is orthogonal to these methods and can be integrated with them to further enhance inference efficiency.
@@ Acknowledgements
We thank Santiago Akle, Vahid Noroozi, Somshubra Majumdar, Jocelyn Huang, Zhiyuan Jerry Lin and NVIDIA Long Context team for helpful discussion and feedback.
@@ B.1 Baseline Comparison
Llama3-8B-Instruct, 1048K Gradient.ai (2024) | 1024 | 32 | 63.70 | -5.32% | 16.9x |
@@
Llama-3.1-70B-Instruct, 128K Meta-AI (2024) | 128 | 16 | 65.29 | -7.47% | 8.7x |
@@ B.2 Hardware for Inference Speed
In Table 1, we use A100 GPUs to run the inference speedup tests. Table 7 describes the number of GPUs and the number of parallel workers used to obtain these numbers for Ring Attention vs Star Attention.
@@
B.3 Prompt Templates
Prompt structure for base models:
Prompt structure for Llama-3 and Llama-3.1 Instruct models:

### v3

@@ 1 Introduction
The importance of long-context capabilities has driven substantial research into addressing the computational challenges of self-attention. Some approaches focus on reducing the need to fully materialize the attention matrix (Milakov & Gimelshein, 2018), leading to blockwise processing techniques (Dao et al., 2022; Dao, 2024; Liu & Abbeel, 2023) and further optimization through distributed computation across multiple compute units (Liu et al., 2024a). While these methods improve training efficiency, autoregressive decoding during inference still requires the model to attend to every previous token, resulting in high computational costs for long-context sequences. Other approaches attempt to optimize inference by segmenting long inputs into chunks, encoding them separately, and attending to these encoded chunks using the user query (Beltagy et al., 2020; Russak et al., 2024; Liao et al., 2024). However, these methods often require fine-tuning the model or introducing additional components that necessitate further training, limiting their out-of-the-box applicability.
We introduce Star Attention111Code: https://github.com/NVIDIA/Star-Attention, a novel algorithm for efficient LLM long-context inference. This method is based on the observation that LLM inference usually has two stages: (1) prompt encoding, where the model processes input and stores KV vectors in the cache and (2) token generation, where model attends to the KV cache and autoregressively generates new tokens while updating the cache with the new KV vectors. In many long-context tasks, the input consists of a long context followed by a short query and a short answer. The information needed for answering the query is often localized within small parts of the context, meaning context tokens need only attend to nearby tokens, while query tokens need to attend to all prior tokens. Based on this observation, Star Attention utilizes a two-phase approach shown in Figure 1:
@@ 2.1 Phase 1: Context Encoding
where each augmented block contains tokens: tokens from the anchor block followed by tokens from the current block (Figure 1). The positional indices of are preserved, ensuring that its tokens retain their original position indices . The augmented blocks are distributed across compute hosts, where each host computes attention over the tokens from its assigned block and generates the corresponding key-value (KV) vectors. While KVs for the anchor block are discarded, the KVs for the current block are retained in the cache.
@@
Model | Seq. Len. | Block Size | Ring-Attn | Star-Attn | |
(K) | (K) | Acc.(%) | Acc. | Speedup | |
16 | 4 | 92.22 | -0.94% | 1.1x | |
32 | 8 | 87.53 | +1.17% | 1.2x | |
64 | 16 | 84.79 | -1.42% | 1.8x | |
Llama-3.1-8B-Instruct (Meta-AI, 2024) | 128 | 32 | 76.31 | -1.90% | 2.7x |
16 | 4 | 95.09 | -2.71% | 1.7x | |
32 | 8 | 94.61 | -2.55% | 2.0x | |
Llama-3.1-70B-Instruct (Meta-AI, 2024) | 64 | 16 | 88.54 | -1.44% | 4.7x |
@@ 3.2 Results
3.3 Comparison with Other Sparse Attention Methods
@@ 3.4 Trade-off between accuracy and speed
Empirically, setting the block size to approximately one-quarter of the total sequence length strikes an effective trade-off between accuracy and speed. For sequence lengths exceeding 128K, we fix the block size at 32K tokens to prioritize inference speed. As shown in Figure 6, this configuration allows Star Attention to achieve substantial speedups over Ring Attention while incurring only modest accuracy degradation. For instance, on the RULER benchmark with Llama-3-8B-Instruct-1048K, Star Attention achieves up to 11 speedup while retaining accuracy comparable to Ring Attention. At 1M tokens, the speedup increases to 16.9 with an accuracy drop of just 5.32%.
@@ 3.5 In-Depth Analysis on RULER Task Categories
To better understand the strengths and limitations of Star Attention, we analyze its performance across different task categories within the RULER benchmark. RULER comprises five categories: Single-NIAH, Multi-NIAH, Multi-Hop Tracing, Aggregation, and Question Answering (QA). Figure 7 reports category-wise accuracy using the Llama-3.1-8B-Instruct model at a sequence length of 32K and a block size of 8K. We observe consistent trends across all sequence lengths, as detailed in Appendix D.
@@ 4 Ablation Study
randomly sampled from [0, current_block) | 96.79 | -2.72% | 97.16 | -1.35% |
same as previous block | 97.35 | -2.16% | 96.80 | -1.71% |
same as first block | 97.61 | -1.90% | 97.54 | -0.96% |
@@
constant token (ex: ‘ ’ or ‘ the’ or ‘.’ ) | 0.00 | -100.00% | 0 | -100.00% |
random tokens | 90.55 | -8.99% | 82.63 | -10.15% |
shuffled first block tokens | 92.96 | -6.57% | 90.76 | -3.26% |
first block tokens | 97.61 | -1.90% | 94.94 | -0.96% |
Previous-block used as anchor | 94.20 | -5.33% | 96.13 | -2.40% |
@@ 5 Related Work
Memory Optimization: Maintaining the KV cache during autoregressive decoding is a major memory bottleneck. KV cache compression (Ge et al., 2024; Munkhdalai et al., 2024; Sun et al., 2024; Liu et al., 2024b; Wu et al., 2024) and low-rank approximation methods (Hu et al., 2022) have been proposed to trade precision for reduced memory. Recent systems explore eviction-based memory management strategies that allow LLMs to operate over virtually infinite contexts (Zhao et al., 2024; Han et al., 2024; Xiao et al., 2024a), often requiring architecture changes or specialized runtime support. Star Attention is orthogonal to these methods and can be integrated with them to further enhance inference efficiency.
@@ Acknowledgements
We thank Kefeng Duan, Santiago Akle, Vahid Noroozi, Somshubra Majumdar, Jocelyn Huang, Zhiyuan Jerry Lin and NVIDIA Long Context team for helpful discussion and feedback.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Star Attention Pseudo-code
Llama3-8B-Instruct, 1048K (Gradient.ai, 2024) | 1024 | 32 | 63.70 | -5.32% | 16.9x |
@@
Llama-3.1-70B-Instruct, 128K (Meta-AI, 2024) | 128 | 16 | 65.29 | -7.47% | 8.7x |
@@ B.2 Hardware for Inference Speed
We use A100 GPUs to run all our inference speedup experiments. Table 7 describes the number of GPUs and the number of parallel workers used to obtain the inference speed numbers for Ring Attention and Star Attention for each sequence length. In all these experiments, the anchor block size in Star Attention was kept same as the context block size.
B.3 Prompt Templates
Prompt template for base models:
Prompt template used for Llama-3 and Llama-3.1 Instruct models: